# Stability of non-supersymmetric vacua from calibrations

### v2

@@
Abstract
@@ 1 Introduction
In this paper, we take a step in that direction: we will extend to supersymmetry-breaking one of the mechanisms that make supersymmetric AdS vacua stable. Powerful positive energy arguments [8, 9, 10, 11, 12] protect the latter from any decay; replicating this success for supersymmetry-breaking is possible in four dimensions [13], but a lot more challenging for compactifications [12]. A much simpler argument was found in [12, Sec. 2] to show protection from decays where the bubble of new vacuum is a D-brane localized at a value of the spacetime radial coordinate. This easier argument is the one we will extend to supersymmetry-breaking vacua.
@@
We have illustrated the idea on several classes of type II vacua. Some of them were already known, and our contribution is their stability analysis; some of them appear to be new, in spite of their simplicity. Here is a list of the classes that we consider:
@@ 3.1 Non-perturbative stability
We begin with a very quick review of vacuum tunneling. This is a quantum effect, and it can be investigated using the path integral approach. After a Wick rotation, the question becomes whether there can exist an instantonic bubble in Euclidean AdS (localized in Euclidean time $\tau$ as well as space) that encloses the new vacuum inside the old. In gravity this was pioneered in [1] for theories with scalars, and extended in [2] for form fields. A simplified setting is when the bubble is a probe brane, whose back-reaction can be neglected. This case was considered in [5, Sec. 4] by solving the equations of motion for the bubble’s shape as a function of $\tau$. In [22, Sec. 4.1] it was shown that this bubble is in fact a half-sphere (bottom part of Fig. 1): in line with earlier expectations from [1], it respects an ${\rm SO}(d)$ subgroup of the ${\rm SO}(d,1)$ isometry group of Euclidean AdS. One can make this symmetry manifest by writing the metric as
@@
So far we have focused on the quantum process that nucleates a bubble of new vacuum. The subsequent evolution can be obtained by an analytic continuation of the above ${\rm SO}(d)$-invariant spherical Euclidean solution, to an ${\rm SO}(d-1,1)$-invariant hyperboloid, representing the evolution of a sphere in AdS [1, 2] (upper part of Fig. 1). As an alternative, one can consider the expansion of branes that are localized along the Poincaré radial coordinate [6, 39]. This leads again to (3.3) [22, Sec. 5.1].
@@ 3.3 Non-supersymmetric solutions
and $\mathrm{e}^{3A-\phi}\text{e}^{-B}\text{Im}\Phi_{2}$ is a proper calibration form. [^6] If there exists at least one generalized submanifold such that the inequality (3.6) is saturated. This eases significantly the stability analysis of $\mathcal{N}=0$ backgrounds satisfying (3.15), since the stability ratio in (3.14) now depends only on the homology class of $\Sigma$. By the RR equations of motion, the integrand in the numerator of (3.14) is closed, so it also only depends on homology classes. In particular, it vanishes for cycles that are homologically trivial.
@@ 4 AdS4 solutions
The reason is that there are only two smooth $\mathrm{QK}_{4}$ manifolds in four dimensions: $\mathbb{CP}^{3}=\mathrm{Tw}(S^{4})$, and the flag manifold $\mathbb{F}(1,2;3)=\mathrm{Tw}(\mathbb{CP}^{2})$ [46]. Both are also homogeneous; for the latter, this property gives rise to additional solutions, so it deserves its own separate treatment in the next subsection. There are additional quaternionic Kähler four-manifolds with orbifold singularities [47, 12.5].
@@ 4.1.2 The solutions
The Ricci tensor for the metric (4.5) can be found in this language in [48, (3.16)]; for $\mathbb{CP}^{3}$, it can also be obtained by considering it as a coset space ${\rm Sp}(2)/{\rm Sp}(1)\times{\rm U}(1)$, using a formula in the next section.
@@
We studied these algebraic equations numerically; our results appear to be in excellent agreement with those in [23, Fig. 2].
@@
As in [55, Sec. 3.2], we take $B=B_{0}+b(j_{\mathrm{F}}-4j_{\mathrm{B}})$; recalling $\mathrm{d}(j_{\mathrm{F}}-4j_{\mathrm{B}})=0$, $b$ is a free parameter. The other free parameters are $\sigma$, $g_{s}$, $L\equiv\sqrt{-3/\Lambda}$. These four are fixed by the four conditions in (4.12). We first compute the periods $n^{0}_{k}\equiv(2\pi)^{1-k}\int_{B_{k}}(\mathrm{e}^{-B_{0}}F)_{k}$; they are of the form $n_{k}^{0}=L^{k-1}/g_{s}\psi_{k}$, where $\psi_{k}$ are multi-valued functions of $\sigma$. The actual flux quanta are then given by a $b$-transform: $n_{0}=n_{0}^{0}$, $n_{2}=n_{2}^{0}-bn_{0}$, and so on. $b$ can be eliminated from the system thanks to two invariants under $b$-transform: $I_{4}=n_{2}^{2}-2n_{0}n_{4}=(n_{2}^{0})^{2}-2n_{0}^{0}n_{4}^{0}$, $I_{6}=n_{2}^{3}+3n_{0}n_{6}-3n_{0}n_{2}n_{4}=(n_{2}^{0})^{3}+3n_{0}^{0}n_{6}^{0}-3n_{0}^{0}n_{2}^{0}n_{4}^{0}$ [55].
@@ 4.2.1 Coset space geometry
Given a fiber bundle $E\buildrel\pi\over{\to}M$, if a form on the total space $E$ is invariant and horizontal (annihilated by both $\iota_{v}$ and $L_{v}=\{\mathrm{d},\iota_{v}\}$, for $v$ any vertical vector field), then it is the pullback under $\pi$ of a form on $M$. In other words, such a form is basically a form on the base.
Consider now the particular case $G\to G/H$. On a group manifold $G$, there are left-invariant one-forms $\lambda^{i}$ obtained by expanding $g^{-1}\mathrm{d}g$ over the generators. In general few (if any) of these are invariant and horizontal under $G\to G/H$. However, it is more common for some higher forms $\frac{1}{k!}\omega_{i_{1}\ldots i_{k}}\lambda^{i_{1}}\wedge\lambda^{i_{k}}$ to be invariant and horizontal. In the context of homogeneous spaces these are often just called invariant. For more details see for example [49, Sec. 3] or [50, Sec. 4.4.2].
@@
Indeed, there are three ways to fiber $\mathbb{F}(1,2;3)$ over $\mathbb{CP}^{2}$ [56, Sec. 5].
@@ 4.2.2 The solutions
To compute the Ricci tensor, we used the following formula [57, (7.33)]:
@@ 4.2.3 Stability: bubbles of bound states
Non-supersymmetric solutions on Kähler–Einstein manifolds $\mathrm{KE}_{6}$ were already considered in [6, 24], and even partially in [25]. This is one of the first examples where bubble stability was addressed [6, Sec. 4.1.2], but only for D2-branes and near the supersymmetric limit in parameter space, for reasons related to the AdS/CFT correspondence. In that limit, a conclusion of instability was reached. Here we will greatly expand that analysis, finding possible islands of stability near the skew-whiffed solutions (already mentioned in Sec. 3.3).
There exist many examples of Kähler–Einstein manifolds. Some famous ones are $\mathbb{CP}^{3}$ and the $d<5$ Fermat equations $\{\sum_{I=1}^{5}Z_{I}^{d}=0\}\subset\mathbb{CP}^{4}$, and the flag manifold ${\mathbb{F}}(1,2;3)$. The problem has famously been reduced [58, 59, 60, 61] to the notion of K-stability. This has been used to great effect in recent years; see e.g. [62, Sec. 6] for a large set of examples.
@@ 4.3.3 Stability: bubbles of bound states
A holomorphic four-cycle $B_{4}$ is simply-connected, by the Lefschetz hyperplane theorem (see e.g. [64, Sec. 1.2]). However, $H^{2}(B_{4},\mathbb{R})$ can be larger than that of the $\mathrm{KE}_{6}$ itself. [^18] For a $B_{4}\subset\mathbb{CP}^{3}$ of degree $d$, the Chern classes are $c_{1}=(4-d)\omega$, $c_{2}=(6-4d+d^{2})\omega^{2}$, with $\omega$ the hyperplane class in $H^{2}(\mathbb{CP}^{3})$. Now the dimension $h_{2}(B_{4})=-2+\chi(B_{4})=-2+\int_{B_{4}}c_{2}=-2+d(6-4d+d^{2})$. On $H^{2}$, the intersection form has signature $(\frac{h_{2}-\sigma}{2},\frac{h_{2}+\sigma}{2})$, $\sigma=\frac{1}{3}\int_{B_{4}}(c_{1}^{2}-2c_{2})=\frac{1}{3}d(4-d^{2})$, but on $H^{1,1}$ it has signature $(1,h^{1,1}-1)$. As a famous example, for $d=4$ we have the quartic K3; it has $h_{2}=22$, signature $(3,19)$ on $H^{2}$, and $(1,19)$ on $H^{1,1}$. Given ${\mathcal{F}}\in H^{2}(B_{4},\mathbb{R})$, let us introduce $k_{1}$, $k_{2}$ such that $\int_{B_{4}}{\mathcal{F}}\wedge{\mathcal{F}}\equiv k_{2}f^{2}\int_{B_{4}}J^{2}$ and $\int_{B_{4}}{\mathcal{F}}\wedge J=k_{1}f\int_{B_{4}}J^{2}$. Due to (3.19b), $f_{\rm ws}$ is $(1,1)$ (and in particular belongs to the Picard lattice $H^{1,1}(B_{4})\cap H^{2}(B_{4},\mathbb{Z})$). By the signature theorem [65, (IV.2.13)], the intersection pairing has signature $(1,h^{1,1}-1)$ on $H^{1,1}$. Applying this to our definitions, one obtains
@@
It might be that this never happens, perhaps along the lines of the Khovanskii–Teissier inequalities [66, Sec. 1.6]. In any case, (4.53) is not an issue when $h^{2}(\mathrm{KE}_{6})=1$, such as for $\mathbb{CP}^{3}$, since in that case ${\cal F}=fJ$ is the only possibility. More generally, (4.53) is less likely to be problematic when $H^{2}(\mathrm{KE}_{6})$ is small. When (4.53) is satisfied, we cannot tell whether there is an actual instability; this would involve solving the six-dimensional case of (3.19a), namely
@@ 5.1 General formalism
The forms $\psi^{a}_{\pm}$ can also be written as bilinears of the spinors on $M_{5}$ obtained by reducing those on $M_{6}$. The Minkowski4 pure spinor equations (which we will see in the next section) can be obtained from (2.8) by sending $\Lambda\to 0$, $L\to\infty$; together with (5.3) they now give [69, (2.15)]
@@ 5.2 Circle bundles over Kähler–Einstein
First we consider $S^{1}$-fibrations over Kähler–Einstein manifolds $\mathrm{KE}_{4}$. These admit the so-called regular Sasaki–Einstein metrics, which are known to admit supersymmetric Freund–Rubin vacua (see [70] and [50, Sec. 7.4] for reviews), but also non-supersymmetric vacua [26, Sec. 3], whose stability we will analyze.
@@ 5.2.1 The fibrations
$\begin{split}0&\to H^{1}(\mathrm{KE}_{4},\mathbb{Z})=0\to H^{1}(\mathrm{SE}_{5},\mathbb{Z})\to H^{0}(\mathrm{KE}_{4},\mathbb{Z})=\mathbb{Z}\\ &\buildrel c_{1}\wedge\over{\to}H^{2}(\mathrm{KE}_{4},\mathbb{Z})\to H^{2}(\mathrm{SE}_{5},\mathbb{Z})\to H^{1}(\mathrm{KE}_{4},\mathbb{Z})=0\\ &\to H^{3}(\mathrm{KE}_{4},\mathbb{Z})\to H^{3}(\mathrm{SE}_{5},\mathbb{Z})\to H^{2}(\mathrm{KE}_{4},\mathbb{Z})\to\ldots\end{split}$ (5.12)
now shows that $H^{k}(\mathrm{SE}_{5},\mathbb{Z})=\{\mathbb{Z},0,\mathbb{Z}^{b-1}\oplus\mathbb{Z}_{m},\mathbb{Z}^{b-1},\mathbb{Z}_{m},\mathbb{Z}\}$, for some $m$ and $b=\mathrm{dim}H^{2}(\mathrm{KE}_{4},\mathbb{R})$ [72, Sec. 4.3]. Using Poincaré duality and the universal coefficient theorem, we also obtain $H_{k}(\mathrm{SE}_{5},\mathbb{Z})=\{\mathbb{Z},\mathbb{Z}_{m},\mathbb{Z}^{b-1},\mathbb{Z}^{b-1}\oplus\mathbb{Z}_{m},0,\mathbb{Z}\}$.

### v3

@@
Stability of non-supersymmetric vacua from calibrations
Vincent Menet and Alessandro Tomasiello
Dipartimento di Matematica, Università di Milano–Bicocca,
Via Cozzi 55, 20126 Milano, Italy
and
INFN, sezione di Milano–Bicocca
vincent.menet, alessandro.tomasiello@unimib.it Abstract
@@ 1 Introduction
In this paper, we take a step in that direction: we will extend to supersymmetry-breaking one of the mechanisms that make supersymmetric AdS vacua stable. Powerful positive energy arguments [72, 29, 37, 41, 30] protect the latter from any decay; replicating this success for supersymmetry-breaking is possible in four dimensions [11], but a lot more challenging for compactifications [30]. A much simpler argument was found in [30, Sec. 2] to show protection from decays where the bubble of new vacuum is a D-brane localized at a value of the spacetime radial coordinate. This easier argument is the one we will extend to supersymmetry-breaking vacua.
@@
We have illustrated the idea on several classes of type II vacua. Some of them were already known, and our contribution is their stability analysis; some of them are new. Here is a list of the classes that we consider:
@@ 3.1 Non-perturbative stability
We begin with a very quick review of vacuum tunneling. This is a quantum effect, and it can be investigated using the path integral approach. After a Wick rotation, the question becomes whether there can exist an instantonic bubble in Euclidean AdS (localized in Euclidean time $\tau$ as well as space) that encloses the new vacuum inside the old. In gravity this was pioneered in [17] for theories with scalars, and extended in [14] for form fields. A simplified setting is when the bubble is a probe brane, whose back-reaction can be neglected. This case was considered in [47, Sec. 4] by solving the equations of motion for the bubble’s shape as a function of $\tau$. In [5, Sec. 4.1] it was shown that this bubble is in fact a half-sphere (bottom part of Fig. 1): in line with earlier expectations from [17], it respects an ${\rm SO}(d)$ subgroup of the ${\rm SO}(d,1)$ isometry group of Euclidean AdS. One can make this symmetry manifest by writing the metric as
@@
So far we have focused on the quantum process that nucleates a bubble of new vacuum. The subsequent evolution can be obtained by an analytic continuation of the above ${\rm SO}(d)$-invariant spherical Euclidean solution, to an ${\rm SO}(d-1,1)$-invariant hyperboloid, representing the evolution of a sphere in AdS [17, 14] (upper part of Fig. 1). As an alternative, one can consider the expansion of branes that are localized along the Poincaré radial coordinate [26, 9]. This leads again to (3.3) [5, Sec. 5.1].
@@ 3.3 Non-supersymmetric solutions
and $\mathrm{e}^{3A-\phi}\text{e}^{-B}\text{Im}\Phi_{2}$ is a proper calibration form. [^6] If there exists at least one generalized submanifold such that the inequality (3.6) is saturated. This eases significantly the stability analysis of $\mathcal{N}=0$ backgrounds satisfying (3.15), since the stability ratio in (3.14) now depends only on the homology class of $\Sigma$. By the RR equations of motion, the integrand in the numerator of (3.14) is closed, so it also only depends on homology classes. [^7] Suppose we know for a basis $\Sigma_{a}$ in homology that $\frac{q(\Sigma_{a})h}{3\tau(\Sigma_{a})}\leq\frac{L\int_{\Sigma_{a}}f}{3\int_{\Sigma_{a}}\psi}\leq 1$, where $f$, $\psi$ are the integrands in (3.14). Then for a general element $\Sigma=s_{a}\Sigma_{a}$, $s_{a}>0$, we have $\frac{q(\Sigma)h}{3\tau(\Sigma)}\leq\frac{Ls_{a}\int_{\Sigma_{a}}f}{3s_{a}\int_{\Sigma_{a}}\psi}\leq 1$. In particular, it vanishes for cycles that are homologically trivial.
@@ 4 AdS4 solutions
The reason is that there are only two smooth $\mathrm{QK}_{4}$ manifolds in four dimensions: $\mathbb{CP}^{3}=\mathrm{Tw}(S^{4})$, and the flag manifold $\mathbb{F}(1,2;3)=\mathrm{Tw}(\mathbb{CP}^{2})$ [35]. Both are also homogeneous; for the latter, this property gives rise to additional solutions, so it deserves its own separate treatment in the next subsection. There are additional quaternionic Kähler four-manifolds with orbifold singularities [13, 12.5].
@@ 4.1.2 The solutions
The Ricci tensor for the metric (4.5) can be found in this language in [67, (3.16)]; for $\mathbb{CP}^{3}$, it can also be obtained by considering it as a coset space ${\rm Sp}(2)/{\rm Sp}(1)\times{\rm U}(1)$, using a formula in the next section.
@@
We studied these algebraic equations numerically; our results appear to be in excellent agreement with those in [38, Fig. 2].
@@
As in [3, Sec. 3.2], we take $B=B_{0}+b(j_{\mathrm{F}}-4j_{\mathrm{B}})$; recalling $\mathrm{d}(j_{\mathrm{F}}-4j_{\mathrm{B}})=0$, $b$ is a free parameter. The other free parameters are $\sigma$, $g_{s}$, $L\equiv\sqrt{-3/\Lambda}$. These four are fixed by the four conditions in (4.12). We first compute the periods $n^{0}_{k}\equiv(2\pi)^{1-k}\int_{B_{k}}(\mathrm{e}^{-B_{0}}F)_{k}$; they are of the form $n_{k}^{0}=L^{k-1}/g_{s}\psi_{k}$, where $\psi_{k}$ are multi-valued functions of $\sigma$. The actual flux quanta are then given by a $b$-transform: $n_{0}=n_{0}^{0}$, $n_{2}=n_{2}^{0}-bn_{0}$, and so on. $b$ can be eliminated from the system thanks to two invariants under $b$-transform: $I_{4}=n_{2}^{2}-2n_{0}n_{4}=(n_{2}^{0})^{2}-2n_{0}^{0}n_{4}^{0}$, $I_{6}=n_{2}^{3}+3n_{0}n_{6}-3n_{0}n_{2}n_{4}=(n_{2}^{0})^{3}+3n_{0}^{0}n_{6}^{0}-3n_{0}^{0}n_{2}^{0}n_{4}^{0}$ [3].
@@ 4.2.1 Coset space geometry
Given a fiber bundle $E\mathrel{\mathop{\kern 0.0pt\to}\limits^{\pi}}M$, if a form on the total space $E$ is invariant and horizontal (annihilated by both $\iota_{v}$ and $L_{v}=\{\mathrm{d},\iota_{v}\}$, for $v$ any vertical vector field), then it is the pullback under $\pi$ of a form on $M$. In other words, such a form is basically a form on the base.
Consider now the particular case $G\to G/H$. On a group manifold $G$, there are left-invariant one-forms $\lambda^{i}$ obtained by expanding $g^{-1}\mathrm{d}g$ over the generators. In general few (if any) of these are invariant and horizontal under $G\to G/H$. However, it is more common for some higher forms $\frac{1}{k!}\omega_{i_{1}\ldots i_{k}}\lambda^{i_{1}}\wedge\lambda^{i_{k}}$ to be invariant and horizontal. In the context of homogeneous spaces these are often just called invariant. For more details see for example [39, Sec. 3] or [68, Sec. 4.4.2].
@@
Indeed, there are three ways to fiber $\mathbb{F}(1,2;3)$ over $\mathbb{CP}^{2}$ [4, Sec. 5].
@@ 4.2.2 The solutions
To compute the Ricci tensor, we used the following formula [10, (7.33)]:
@@ 4.2.3 Stability: bubbles of bound states
Non-supersymmetric solutions on Kähler–Einstein manifolds $\mathrm{KE}_{6}$ were already considered in [26, 45], and even partially in [60]. This is one of the first examples where bubble stability was addressed [26, Sec. 4.1.2], but only for D2-branes and near the supersymmetric limit in parameter space, for reasons related to the AdS/CFT correspondence. In that limit, a conclusion of instability was reached. Here we will greatly expand that analysis, finding possible islands of stability near the skew-whiffed solutions (already mentioned in Sec. 3.3).
There exist many examples of Kähler–Einstein manifolds. Some famous ones are $\mathbb{CP}^{3}$ and the $d<5$ Fermat equations $\{\sum_{I=1}^{5}Z_{I}^{d}=0\}\subset\mathbb{CP}^{4}$, and the flag manifold ${\mathbb{F}}(1,2;3)$. The problem has famously been reduced [75, 65, 16, 66] to the notion of K-stability. This has been used to great effect in recent years; see e.g. [7, Sec. 6] for a large set of examples.
@@ 4.3.3 Stability: bubbles of bound states
A holomorphic four-cycle $B_{4}$ is simply-connected, by the Lefschetz hyperplane theorem (see e.g. [32, Sec. 1.2]). However, $H^{2}(B_{4},\mathbb{R})$ can be larger than that of the $\mathrm{KE}_{6}$ itself. [^19] For a $B_{4}\subset\mathbb{CP}^{3}$ of degree $d$, the Chern classes are $c_{1}=(4-d)\omega$, $c_{2}=(6-4d+d^{2})\omega^{2}$, with $\omega$ the hyperplane class in $H^{2}(\mathbb{CP}^{3})$. Now the dimension $h_{2}(B_{4})=-2+\chi(B_{4})=-2+\int_{B_{4}}c_{2}=-2+d(6-4d+d^{2})$. On $H^{2}$, the intersection form has signature $(\frac{h_{2}-\sigma}{2},\frac{h_{2}+\sigma}{2})$, $\sigma=\frac{1}{3}\int_{B_{4}}(c_{1}^{2}-2c_{2})=\frac{1}{3}d(4-d^{2})$, but on $H^{1,1}$ it has signature $(1,h^{1,1}-1)$. As a famous example, for $d=4$ we have the quartic K3; it has $h_{2}=22$, signature $(3,19)$ on $H^{2}$, and $(1,19)$ on $H^{1,1}$. Given ${\mathcal{F}}\in H^{2}(B_{4},\mathbb{R})$, let us introduce $k_{1}$, $k_{2}$ such that $\int_{B_{4}}{\mathcal{F}}\wedge{\mathcal{F}}\equiv k_{2}f^{2}\int_{B_{4}}J^{2}$ and $\int_{B_{4}}{\mathcal{F}}\wedge J=k_{1}f\int_{B_{4}}J^{2}$. Due to (3.19b), $f_{\rm ws}$ is $(1,1)$ (and in particular belongs to the Picard lattice $H^{1,1}(B_{4})\cap H^{2}(B_{4},\mathbb{Z})$). By the signature theorem [8, (IV.2.13)], the intersection pairing has signature $(1,h^{1,1}-1)$ on $H^{1,1}$. Applying this to our definitions, one obtains
@@
It might be that this never happens, perhaps along the lines of the Khovanskii–Teissier inequalities [42, Sec. 1.6]. In any case, (4.53) is not an issue when $h^{2}(\mathrm{KE}_{6})=1$, such as for $\mathbb{CP}^{3}$, since in that case ${\cal F}=fJ$ is the only possibility. More generally, (4.53) is less likely to be problematic when $H^{2}(\mathrm{KE}_{6})$ is small. When (4.53) is satisfied, we cannot tell whether there is an actual instability; this would involve solving the six-dimensional case of (3.19a), namely
@@ 5.1 General formalism
The forms $\psi^{a}_{\pm}$ can also be written as bilinears of the spinors on $M_{5}$ obtained by reducing those on $M_{6}$. The Minkowski4 pure spinor equations (which we will see in the next section) can be obtained from (2.8) by sending $\Lambda\to 0$, $L\to\infty$; together with (5.3) they now give [6, (2.15)]
@@ 5.2 Circle bundles over Kähler–Einstein
First we consider $S^{1}$-fibrations over Kähler–Einstein manifolds $\mathrm{KE}_{4}$. These admit the so-called regular Sasaki–Einstein metrics, which are known to admit supersymmetric Freund–Rubin vacua (see [62] and [68, Sec. 7.4] for reviews), but also non-supersymmetric vacua [59, Sec. 3], whose stability we will analyze.
@@ 5.2.1 The fibrations
$\begin{split}0&\to H^{1}(\mathrm{KE}_{4},\mathbb{Z})=0\to H^{1}(\mathrm{SE}_{5},\mathbb{Z})\to H^{0}(\mathrm{KE}_{4},\mathbb{Z})=\mathbb{Z}\\ &\mathrel{\mathop{\kern 0.0pt\to}\limits^{c_{1}\wedge}}H^{2}(\mathrm{KE}_{4},\mathbb{Z})\to H^{2}(\mathrm{SE}_{5},\mathbb{Z})\to H^{1}(\mathrm{KE}_{4},\mathbb{Z})=0\\ &\to H^{3}(\mathrm{KE}_{4},\mathbb{Z})\to H^{3}(\mathrm{SE}_{5},\mathbb{Z})\to H^{2}(\mathrm{KE}_{4},\mathbb{Z})\to\ldots\end{split}$ (5.12)
now shows that $H^{k}(\mathrm{SE}_{5},\mathbb{Z})=\{\mathbb{Z},0,\mathbb{Z}^{b-1}\oplus\mathbb{Z}_{m},\mathbb{Z}^{b-1},\mathbb{Z}_{m},\mathbb{Z}\}$, for some $m$ and $b=\mathrm{dim}H^{2}(\mathrm{KE}_{4},\mathbb{R})$ [55, Sec. 4.3]. Using Poincaré duality and the universal coefficient theorem, we also obtain $H_{k}(\mathrm{SE}_{5},\mathbb{Z})=\{\mathbb{Z},\mathbb{Z}_{m},\mathbb{Z}^{b-1},\mathbb{Z}^{b-1}\oplus\mathbb{Z}_{m},0,\mathbb{Z}\}$.